# Constraints on Dark Matter with Future MeV Gamma-Ray Telescopes

### v2

@@ IV.4 Comparison with the literature and cross-checks
We note briefly that there are discrepancies between our projected constraints for GECCO and those from [5]; our results differ by various constant factors (typically $\mathcal{O}(1)$ and in all cases less than one order of magnitude).
To derive a cross-check on our results, let us consider a simplified version of our analysis. Since the energy resolution of GECCO is quite narrow, the energy spectrum of photons can be approximated by a delta function, for purposes of estimating the background to a line signal. In this limit, the ratio of the Galactic Center constraints to the M31 constraints should be approximately equal to the ratio of the decay factors divided by the square root of the ratio of the background models, evaluated at the desired mass point. We also find in this limit that, for the method we used,
@@
where $\langle\sigma v\rangle$ is the annihilation constraint calculated at $m_{\chi}=m_{a}$ with an observation angle of $\Omega_{a}$ and a $J$-factor of $J_{a}$, $\tau$ is the decay constraint calculated for the same target and integration time at $m_{\chi}=m_{d}=2m_{a}$ with an observation angle of $\Omega_{d}$ and a $J$-factor of $J_{d}$, and $f_{\chi}=1$ if the particles are self-conjugate and $f_{\chi}=2$ otherwise. We find that our calculations for all of the constraints for the photon annihilation and decay modes, in cases where the telescope energy resolution is good and these approximations should be valid, are consistent with these relations.
We also compared our results for AdEPT with Draco observations to those of Ref. [15]. They assume a much longer run time (5 years vs $10^{6}$ seconds) and also take an energy-independent value for the angular resolution that is smaller than our prescription across much of the energy range, allowing them to set a small ROI for their dwarf analysis which consequently reduces background. The $J$- and $D$-factors they assume for Draco are also modestly different from ours. Once these effects are taken into account, our forecast limits appear roughly consistent for the decay case and a factor of a few weaker for the annihilation case; we do employ a different statistical method so this may be responsible for at least some of the difference.
@@ V Conclusions
All of the instruments we considered would markedly improve our current constraints for dark matter decay or annihilation directly to photons in their respective energy ranges when pointed at the Galactic Center. MAST, e-ASTROGAM, AMEGO, PANGU, and AdEPT would all also improve our current constraints for decay or annihilation channels involving neutral pions when pointed at the Galactic Center. However, since all of the instruments we considered are photon instruments, they are not able to place constraints that are as strong onto lepton decay or annihilation modes, as we have focused only on photons produced directly in the annihilation/decay (via radiative corrections), rather than lower-energy secondary photons that would require modeling of the charged particles’ propagation. That being said, for lower-energy dark matter, GECCO, GRAMS, e-ASTROGAM, and AMEGO are all forecast to be able to place competitive constraints on decay or annihilation to electrons. Additionally, MAST is forecast to be able to improve constraints on annihilation to leptons if pointed at the Galactic Center given a long enough integration time. All the constraints shown were for an integration time of $10^{6}$ seconds, and it is reasonable to extend this integration time by a couple of orders of magnitude in the event that one of these instruments is mounted on a satellite, given that none of the instruments are pointed.
Acknowledgements
The authors thank Stefano Profumo, Tsuguo Aramaki, and Regina Caputo for helpful discussions, and Dave Tucker-Smith for pointing out a plotting error in v1 of this work. K.E.O was supported by the MIT Undergraduate Research Opportunities Program, by the Mariana Polonsky Slocum (1955) Memorial Fund, and by the DeFlorez Endowment Fund. T.R.S. was supported in part by a Guggenheim Fellowship; the Edward, Frances, and Shirley B. Daniels Fellowship of the Harvard Radcliffe Institute; the Bershadsky Distinguished Fellowship of the Harvard Physics Department; and the Simons Foundation (Grant Number 929255, T.R.S). T.R.S. thanks the Kavli Institute for Theoretical Physics (KITP), the Aspen Center for Physics, and the Mainz Institute for Theoretical Physics for their hospitality during the completion of this work; this research was supported in part by grant no. NSF PHY-2309135 to KITP, and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. This work was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567.
@@ Appendix A Results for Draco and M31
Figure 7: The constraints each instrument, pointed at Draco, can place on dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in Section IV.
Figure 8: The constraints each instrument, pointed at Draco, can place on dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in Section IV.
In Figures 7 and 8, we display our results for each instrument and decay/annihilation mode for observation of Draco. We note that these constraints are proportional to the decay factor, so the M31 constraints are simply a rescaling of the Draco constraints since we used the same background model for both targets.
@@ Appendix B Atmospheric backgrounds
Figure 9: The constraints that the balloon version of GRAMS (blue) and e-ASTROGAM (orange), pointed at the Galactic Center, can place on dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$, taking atmospheric backgrounds into consideration (dashed lines). These are compared against the results we obtained in Section IV without accounting for atmospheric backgrounds (solid lines).
Figure 10: The constraints that the balloon version of GRAMS (blue) and e-ASTROGAM (orange), pointed at the Galactic Center, can place on dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$, taking atmospheric backgrounds into consideration (dashed lines). These are compared against the results we obtained in Section IV without accounting for atmospheric backgrounds (solid lines).
We now explore how the inclusion of atmospheric backgrounds affects our results. Since there do not yet exist atmospheric background estimates for most of the instruments we consider in this work, we focus on e-ASTROGAM and the balloon version of GRAMS in this section as examples of time-projection chamber and silicon-tracker technology respectively. An atmospheric background estimate for GRAMS is given in Figure 6 of Ref. [21], and an atmospheric background estimate for e-ASTROGAM is given in Figure 18 of Ref. [37]. In both cases, we assume the atmospheric background is isotropic and well-measured (i.e. we do not include any free parameters associated with the atmospheric background in our Fisher analysis). The effect of including the atmospheric backgrounds is plotted in Figures 9 and 10. Overall, we find that accounting for atmospheric backgrounds changes our results by a factor of a few in the case of e-ASTROGAM and less than an order of magnitude in the case of GRAMS. Strategies to further reduce these backgrounds, or to effectively subtract them (e.g. by cuts or by direct measurement of the background in a region where it is bright) could mitigate this modest loss of sensitivity.

### v3

@@ IV.4 Comparison with the literature and cross-checks
We note briefly that there are discrepancies between our projected constraints for GECCO and those from [5]; our results differ by various constant factors [typically $\mathcal{O}(1)$ and in all cases less than one order of magnitude].
To derive a cross-check on our results, let us consider a simplified version of our analysis. Since the energy resolution of GECCO is quite narrow, the energy spectrum of photons can be approximated by a $\delta$ function, for purposes of estimating the background to a line signal. In this limit, the ratio of the Galactic Center constraints to the M31 constraints should be approximately equal to the ratio of the decay factors divided by the square root of the ratio of the background models, evaluated at the desired mass point. We also find in this limit that, for the method we used,
@@
where $\langle\sigma v\rangle$ is the annihilation constraint calculated at $m_{\chi}=m_{a}$ with an observation angle of $\Omega_{a}$ and a $J$ factor of $J_{a}$, $\tau$ is the decay constraint calculated for the same target and integration time at $m_{\chi}=m_{d}=2m_{a}$ with an observation angle of $\Omega_{d}$ and a $J$ factor of $J_{d}$, and $f_{\chi}=1$ if the particles are self-conjugate and $f_{\chi}=2$ otherwise. We find that our calculations for all of the constraints for the photon annihilation and decay modes, in cases where the telescope energy resolution is good and these approximations should be valid, are consistent with these relations.
We also compared our results for AdEPT with Draco observations to those of Ref. [16]. They assume a much longer run time (5 yr vs $10^{6}$ s) and also take an energy-independent value for the angular resolution that is smaller than our prescription across much of the energy range, allowing them to set a small ROI for their dwarf analysis which consequently reduces background. The $J$ and $D$ factors they assume for Draco are also modestly different from ours. Once these effects are taken into account, our forecast limits appear roughly consistent for the decay case and a factor of a few weaker for the annihilation case; we do employ a different statistical method so this may be responsible for at least some of the difference.
@@ V Conclusions
All of the instruments we considered would markedly improve our current constraints for dark matter decay or annihilation directly to photons in their respective energy ranges when pointed at the Galactic Center. MAST, e-ASTROGAM, AMEGO, PANGU, and AdEPT would all also improve our current constraints for decay or annihilation channels involving neutral pions when pointed at the Galactic Center. However, since all of the instruments we considered are photon instruments, they are not able to place constraints that are as strong onto lepton decay or annihilation modes, as we have focused only on photons produced directly in the annihilation/decay (via radiative corrections), rather than lower-energy secondary photons that would require modeling of the charged particles’ propagation. That being said, for lower-energy dark matter, GECCO, GRAMS, e-ASTROGAM, and AMEGO are all forecast to be able to place competitive constraints on decay or annihilation to electrons. Additionally, MAST is forecast to be able to improve constraints on annihilation to leptons if pointed at the Galactic Center given a long enough integration time. All the constraints shown were for an integration time of $10^{6}$ s, and it is reasonable to extend this integration time by a couple of orders of magnitude in the event that one of these instruments is mounted on a satellite, given that none of the instruments are pointed.
Acknowledgments
The authors thank Stefano Profumo, Tsuguo Aramaki, and Regina Caputo for helpful discussions, and Dave Tucker-Smith for pointing out a plotting error in an earlier version of this work. K.E.O was supported by the MIT Undergraduate Research Opportunities Program, by the Mariana Polonsky Slocum (1955) Memorial Fund, and by the DeFlorez Endowment Fund. T.R.S. was supported in part by a Guggenheim Fellowship; the Edward, Frances, and Shirley B. Daniels Fellowship of the Harvard Radcliffe Institute; the Bershadsky Distinguished Fellowship of the Harvard Physics Department; and the Simons Foundation (Grant No. 929255). T.R.S. thanks the Kavli Institute for Theoretical Physics (KITP), the Aspen Center for Physics, and the Mainz Institute for Theoretical Physics for their hospitality during the completion of this work; this research was supported in part by Grant No. NSF PHY-2309135 to KITP and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation Grant No. PHY-2210452. This work was supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under Contract No. DE-SC0012567.
@@ Appendix A Results for Draco and M31
Figure 7: The constraints each instrument, pointed at Draco, can place on dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in Sec. IV.
Figure 8: The constraints each instrument, pointed at Draco, can place on dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$. These are compared against the current constraints described in Sec. IV.
In Figs. 7 and 8, we display our results for each instrument and decay/annihilation mode for observation of Draco. We note that these constraints are proportional to the decay factor, so the M31 constraints are simply a rescaling of the Draco constraints since we used the same background model for both targets.
@@ Appendix B Atmospheric backgrounds
Figure 9: The constraints that the balloon version of GRAMS (blue) and e-ASTROGAM (orange), pointed at the Galactic Center, can place on dark matter decay and annihilation into (from top to bottom) two photons, a photon and a neutral pion, and two neutral pions, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$, taking atmospheric backgrounds into consideration (dashed lines). These are compared against the results we obtained in Sec. IV without accounting for atmospheric backgrounds (solid lines).
Figure 10: The constraints that the balloon version of GRAMS (blue) and e-ASTROGAM (orange), pointed at the Galactic Center, can place on dark matter decay and annihilation into the final states (from top to bottom) $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, and $\pi^{+}\pi^{-}$, assuming an observation time of $T_{\mathrm{obs}}=10^{6}\mathrm{\ s}$, taking atmospheric backgrounds into consideration (dashed lines). These are compared against the results we obtained in Sec. IV without accounting for atmospheric backgrounds (solid lines).
We now explore how the inclusion of atmospheric backgrounds affects our results. Since there do not yet exist atmospheric background estimates for most of the instruments we consider in this work, we focus on e-ASTROGAM and the balloon version of GRAMS in this section as examples of time-projection chamber and silicon-tracker technology respectively. An atmospheric background estimate for GRAMS is given in Fig. 6 of Ref. [22], and an atmospheric background estimate for e-ASTROGAM is given in Fig. 18 of Ref. [39]. In both cases, we assume the atmospheric background is isotropic and well measured (i.e. we do not include any free parameters associated with the atmospheric background in our Fisher analysis). The effect of including the atmospheric backgrounds is plotted in Figs. 9 and 10. Overall, we find that accounting for atmospheric backgrounds changes our results by a factor of a few in the case of e-ASTROGAM and less than an order of magnitude in the case of GRAMS. Strategies to further reduce these backgrounds, or to effectively subtract them (e.g. by cuts or by direct measurement of the background in a region where it is bright) could mitigate this modest loss of sensitivity.